OpenStudy (anonymous):

Find the y-intercept of the graph of the equation. 2x+5y=30

OpenStudy (anonymous):

@happytales

OpenStudy (anonymous):

DUDE!!! 6... just change the equation to match y=mx+b. It should be simple at this point.

OpenStudy (anonymous):

yea but I need to show the work that's why I ask.

OpenStudy (anonymous):

I don't know how to show the work ?

OpenStudy (anonymous):

2x+5y=30 Subtract 2x 5y=2x+30 Divide: y=2/5x+6